OpenStudy (anonymous):

Please help. seven-fourths decreased by a number is 2.

OpenStudy (phi):

they want you to figure out what x is in \[ \frac{7}{4}- x = 2\]

OpenStudy (anonymous):

tell me if I'm on the right track please. 1.75=1 3/4-n=2

OpenStudy (texaschic101):

well.....you can either add x to both sides and then subtract 2 from both sides or you can subtract 7/4 from both sides, then divide by -1. It is not necessary to change 7/4 into a mixed number.

OpenStudy (whpalmer4):

Easiest way might be to clear the fractions by multiplying everything by 4: \[4*\frac{7}{4}-4*x = 4*2\]\[7 - 4x = 8\]

OpenStudy (texaschic101):

or you can multiply the entire equation by 4

OpenStudy (whpalmer4):

now subtract 7 from both sides and divide both sides by -4 and you'll have your answer. Notice that the only way to "decrease" 7/4 and get 2 is by subtracting a negative number...

OpenStudy (anonymous):

-1/4

OpenStudy (whpalmer4):

That's correct. \[\frac{7}{4} -(-\frac{1}{4}) = \frac{7-(-1)}{4} = \frac{7+1}{4} = \frac{8}{4} = 2\]